# The Sound of Silence in Social Networks

### v1

@@ 2.3.2 Memory-based Update
Let $M=(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an SO model where $G=(A,E,I)$ is an $n$-agent influence graph. Then $M$ is said to be a memory-based SO model ($\mbox{SOM}^{+}$) if $\mathbf{S}^{0}=\mathbf{1}_{n}$ and for each $i\in A$ and $t\in\mathbb{N}$, $\mu_{G}(\mathbf{B}^{t},\mathbf{S}^{t},t)=(\mathbf{B}^{t+1},\mathbf{S}^{t+1})$ where $\mathbf{B}^{t+1}_{i}$ and $\mathbf{S}^{t+1}_{i}$ are determined by Eq .4 and Eq .5, respectively.
@@
The main difference between $\mbox{SOM}^{+}$ and the DeGroot model (and $\mbox{SOM}^{-}$) is that the latter is a Markovian process: The next state depends on the current state but not past states. In fact, much of the tractability of the DeGroot model derives from its connection to Markov chains. Nevertheless, the next state in $\mbox{SOM}^{+}$ does not depend on the entire state history but just on the most recent public opinions. In the next sections, we will see the impact of this minimal amount of memory on opinion evolution.
@@ 3 Results on Consensus
Consensus is a central problem in social learning models. Often, an inability to reach a consensus serves as an indicator of polarization within a social network. In our model, consensus represents convergence to the same opinion value over time.
@@
In this section, we explore the different results on consensus for both our models on two different graph topologies clique and strongly connected. We find that consensus can only be guaranteed for the $\mbox{SOM}^{-}$ model on clique graphs. In the remaining 3 cases, consensus cannot be guaranteed, leading to potential dissensus due to the existence of perpetual silence for $\mbox{SOM}^{-}$ and public opinions for $\mbox{SOM}^{+}$. We show the counterexamples via simulations.
@@ 3.1 $\mbox{SOM}^{-}$ Properties
A key property of the $\mbox{SOM}^{-}$ model is that if all agents become silent at time $t$, they will all speak up at the very next round $t+1$. The following lemma formalizes this property:
Lemma 1
@@
As a result of Lem. 1, it follows that in $\mbox{SOM}^{-}$ we cannot have all agents silent forever.
Corollary 1
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},$ $\mu_{G})$ be an $\mbox{SOM}^{-}$ model with $G=(A,E,I)$. For $\text{For any}\ t\in\mathbb{N}$,there $\ \text{there exist}\ i\in A$ such that $\mathbf{S}^{t}_{i}=1$ or $\mathbf{S}^{t+1}_{i}=1$.
We shall now prove that the sequences of maximum and minimum opinion values, $\{\max(\mathbf{B}^{t})\}_{t\in\mathbb{N}}$ and $\{\min(\mathbf{B}^{t})\}_{t\in\mathbb{N}}$, are (bounded) monotonically non-increasing and non-decreasing, respectively, so they must converge to some opinion values, say $U$ and $L$ with $L\leq U$.
Lemma 2 (Bounds of Opinion Evolution)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an $\mbox{SOM}^{-}$ model with $G=(A,E,I)$. For $\text{For any}\ t\in\mathbb{N}$, $min(\mathbf{B}^{t})\leq\mathbf{B}^{t+1}_{i}\leq max(\mathbf{B}^{t})$ for all $i\in A$.
@@
$\mathbf{B}^{t+1}_{i}=\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}^{t}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot(\mathbf{B}^{t}_{j}-\mathbf{B}^{t}_{i})=\sum_{j\in N_{i}^{t}\cup\{i\}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}$
Now, as $\mathbf{B}^{t}_{j}\leq max(\mathbf{B}^{t})$, $\mathbf{S}^{t}_{j}\in\{0,1\}$, and $\sum_{j\in N_{i}^{t}\cup\{i\}}I_{ji}$ $=1$ by Definition 1, we can conclude that:
$\mathbf{B}^{t+1}_{i}\leq\sum_{j\in N_{i}^{t}\cup\{i\}}I_{ji}\cdot 1\cdot\mathbf{B}^{t}_{j}\leq\sum_{j\in N_{i}^{t}\cup\{i\}}I_{ji}\cdot 1\cdot max(\mathbf{B}^{t})=1\cdot 1\cdot max(\mathbf{B}^{t})$
As wanted. The proof that $min(\mathbf{B}^{t})\leq\mathbf{B}^{t+1}_{i}$ is similar.
Corollary 2 (Monotonicity of extremes)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an $\mbox{SOM}^{-}$ model with $G=(A,E,I)$. For all $t\in\mathbb{N},$ $\ max(\mathbf{B}^{t+1})$ $\leq max(\mathbf{B}^{t})$ and $\ min(\mathbf{B}^{t+1})\geq min(\mathbf{B}^{t})$
Monotonicity and boundedness of extremes, together with the Monotonic Convergence Theorem [24], lead us to the existence of limits for the opinion values of extreme agents.
Theorem 3.1 (Limits of extremes)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an $\mbox{SOM}^{-}$ model with $G=(A,E,I)$. There exist $\ U,\,L\in[0,1]$ such that $\ \lim_{t\to\infty}\{max(\mathbf{B}^{t})\}=U$ and $\ \lim_{t\to\infty}\{min(\mathbf{B}^{t})\}$ $=L$.
Notice that if the limits for the opinion values of extreme agents are the same (i.e., $U=L$) and by the squeeze theorem [24], we know the following:
$\lim_{t\to\infty}max(\mathbf{B}^{t})=\lim_{t\to\infty}min(\mathbf{B}^{t})=\lim_{t\to\infty}\mathbf{B}^{t}_{i}\;\;\forall i\in A$
Hence, if $U=L$, the model converges to consensus.
3.2 Consensus in $\mbox{SOM}^{-}$  Cliques
In this section we show that consensus is guaranteed for $\mbox{SOM}^{-}$ models whose influence graphs are cliques with at least three agents.
The proof strategy is based on the following key observations: 1. From any time $t$ onwards, there will always be non-silent agents with an opinion value greater or equal to $U$ (Lem. 3). 2. The maximum and minimum opinion values, which are monotonically non-increasing and non-decreasing respectively, must converge to some value, say $U$ and $L$ respectively, with $L\leq U$ (Th. 3.1). 3. As the graph is a clique, the high opinion non-silent agents (i.e., those non-silent agents with opinion value $\geq$ $U$) will influence all other graph agents infinitely often (Lem. 3. 4. Consider any clique with size $n\geq 3$, if $L<U$, the difference between the maximum and minimum opinions will be reduced by a constant factor infinitely often, such that eventually the decrement would be greater than the difference between the maximum and minimum opinions; a contradiction. Hence, $L=U$; which guarantees convergence to consensus.
Remember that in an $\mbox{SOM}^{-}$ model we cannot have all the agents becoming silent forever (Cor. 1). Therefore, at least one agent will be non-silent in the future from any time; specifically, we can also show that from any time onwards, in a clique, at least an agent with an opinion value greater or equal to $U$ will be non-silent infinitely often.
Lemma 3 (silence and maximal opinions)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be a $\mbox{SOM}^{-}$ model with $G=(A,E,I)$ where $G$ is a clique. For $\text{For any}\ t\in\mathbb{N}$, there $\text{there exists}\ k\in\mathbb{N}$ and $\ i\in A$, such that  $\mathbf{B}^{t+k}_{i}\geq U$ and  $\mathbf{S}^{t+k}_{i}=1$.

### v2

@@ 2.3.1 Memoryless Update
Let $M=(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an SO model where $G=(A,E,I)$ is an $n$-agent influence graph. Then $M$ is said to be an SO memory-based ($\mbox{SOM}^{+}$) model if $\mathbf{S}^{0}=\mathbf{1}_{n}$ and for each $i\in A$ and $t\in\mathbb{N}$, $\mu_{G}(\mathbf{B}^{t},\mathbf{S}^{t},t)=(\mathbf{B}^{t+1},\mathbf{S}^{t+1})$ where $\mathbf{B}^{t+1}_{i}$ and $\mathbf{S}^{t+1}_{i}$ are determined by Eq. 5 and Eq. 6, resp.
@@
The main difference between $\mbox{SOM}^{+}$ and the DeGroot and $\mbox{SOM}^{-}$ models is that the latter two are Markovian processes. The next state depends on the current state but not past states. In fact, much of the tractability of DeGroot models derives from its connection to Markov chains. Nevertheless, the next state in $\mbox{SOM}^{+}$ models does not depend on the entire state history but just on the most recent public opinions. In the next sections, we will see the impact of this limited amount of memory on opinion evolution.
@@ 3 Results on Consensus
Consensus is a central problem in social learning models. Often, an inability to reach a consensus is a sign of polarization. In the DeGroot framework, consensus represents convergence to the same opinion value over time.
@@
In this section, we explore consensus for both types of models on two different graph topologies: clique and strongly connected graphs. We show that consensus can only be guaranteed in arbitrary $\mbox{SOM}^{-}$ models on clique graphs. In all other cases, consensus cannot be guaranteed due to the existence of perpetual silence in the case of $\mbox{SOM}^{-}$ models and public opinions in the case for $\mbox{SOM}^{+}$ models.
Due to space limitations, in this section we provide only sketches of the proofs for the lemmas and theorems; the complete proofs can be found in the Appendix.
@@ 3.1 $\mbox{SOM}^{-}$ Properties
A key property of $\mbox{SOM}^{-}$ models is that if all agents become silent at time $t$, they will all speak up at the very next round $t+1$. The following lemma formalizes this property:
Lemma 1 ()
@@
From Lem. 1, in $\mbox{SOM}^{-}$ models we cannot have all agents silent forever.
Corollary 1
[…]
We will now prove that the sequences of maximum and minimum opinion values, $\{\max(\mathbf{B}^{t})\}_{t\in\mathbb{N}}$ and $\{\min(\mathbf{B}^{t})\}_{t\in\mathbb{N}}$, are (bounded) monotonically non-increasing and non-decreasing, respectively, so they must converge to some opinion value, say $U$ and $L$ with $L\leq U$. In Section 3.2, we will prove that, under certain conditions, $U=L$, which implies that the model converges to consensus.
Lemma 2 (Opinion Bounds)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an $\mbox{SOM}^{-}$ model with $G=(A,E,I)$. $\text{For any}\ t\in\mathbb{N}$, $min(\mathbf{B}^{t})\leq\mathbf{B}^{t+1}_{i}\leq max(\mathbf{B}^{t})$ for all $i\in A$.
Corollary 2 (Monotonicity of Extremes)
[…]
The monotonicity (and boundedness) of extremes and the Monotonic Convergence Theorem [27], lead us to the existence of limits for the opinion values of extreme agents.
Theorem 3.1 (Limits of Extremes)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an $\mbox{SOM}^{-}$ model with $G=(A,E,I)$. There must exist $U,\,L\in[0,1]$ such that $\lim_{t\to\infty}\{max(\mathbf{B}^{t})\}=U$ and $\lim_{t\to\infty}\{min(\mathbf{B}^{t})\}$ =L=L.
Henceforth, $U$ and $L$ refer to the limit values from Th. 3.1. Notice that if the limits for the opinion values of extreme agents are the same (i.e., $U=L$) and by the squeeze theorem [27], we can conclude that all the agents converge to consensus.
3.2 Consensus in $\mbox{SOM}^{-}$  Cliques
In this section, we show that consensus is guaranteed for $\mbox{SOM}^{-}$ models whose influence graphs are cliques with at least three agents.
We consider the minimum influence of the underlying (clique) graph $G=(A,E,I)$, defined as the constant $I_{min}=$ $\min_{(i,j)\in E}$ $I(i,j)$  [^3] Notice that $I_{min}>0$ since we are assuming that the underlying graph is a clique. and the difference between the maximum and minimum opinion values at time $t$, defined as $R_{t}=max(\mathbf{B}^{t})-\min(\mathbf{B}^{t})$, notice that $R_{t}\in[0,1]$ at any time $t$.
The proof strategy is based on the following key observations: (a) From any time $t$ onwards, there will always be non-silent agents (Cor. 1). (b) The maximum and minimum opinion values, which are monotonically non-increasing and non-decreasing respectively, must converge to some value, say $U$ and $L$ respectively, with $L\leq U$ (Th. 3.1). (c) As the graph is a clique, the non-silent agents will influence, through the update function, all other graph agents infinitely often (Cor. 1). In each such update, the span between the maximum and minimum opinions is multiplied by at most the constant $1-I_{min}<1$ (Lem. 3). Since this contraction recurs infinitely often, the span tends to zero, which forces all opinions to converge to a common value (Th. 3.2).
To prove consensus (i.e., $U=L$), we now demonstrate how the difference between the maximum and minimum opinion values is reduced over different time units.
Lemma 3 (Convergent Extremes Differences)
Let $(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ be an $\mbox{SOM}^{-}$ model with an $n$-agent influence graph $G=(A,E,I)$ where $G$ is a clique with $n\geq 3$. For all $m\in\mathbb{N}$, there exists $t\in\mathbb{N}$ such that $R_{t}\leq R_{0}\cdot(1-I_{min})^{m}$.
As the difference between the maximum and minimum opinions reduces over time, approaching zero, we can now state our consensus result for cliques with at least three agents.
Theorem 3.2 (Consensus in $\mbox{SOM}^{-}$ Cliques)
Remark 3
Notice that for $2$-agent cliques ($n=2$) consensus is not guaranteed; let us consider a clique with two agents: Agent $1$ and Agent $2$, where the opinions are $\mathbf{B}^{0}_{1}=1$ and $\mathbf{B}^{0}_{2}=0$, the influences are $I(1,2)=I(2,1)=1$, the tolerance radii are $\tau_{1}=\tau_{2}=1$ and the majority thresholds are $\mathcal{M}_{1}=\mathcal{M}_{2}=0.5$. In this case, the opinion evolution of agent $1$, starting with opinion $1$, and agent $2$, starting with opinion 0, always alternates between the values $1$ and 0. Figure 1 illustrates the opinion evolution of this clique.
(a) Influence graph
(b) State evolution
Figure 1: 2-agent clique opinion evolution; $\mathbf{B}^{0}=(1,0)$, $\tau_{i}=1$, $\mathcal{M}_{i}=0.5,i\in\{1,2\}$.
3.3 Dissensus in $\mbox{SOM}^{-}$ Models
In strongly connected aperiodic graphs, the $\mbox{SOM}^{-}$ models differs from DeGroot models by no longer guaranteeing consensus. Agents can enter a state of perpetual silence, effectively disrupting opinion propagation as if severing connections from the graph. This phenomenon is particularly critical when silent agents form bridges between connected components. Their opinions, influenced by opposing connected components, may remain below their majo threshold indefinitely. As a result, they prevent opinion exchange between components, obstructing the possibility of achieving consensus. The following example illustrates this scenario by showing the agents state evolution and influence graph.
Remark 4
For visual clarity, self-influences are omitted. As a result, the visible incoming influences for each agent may not sum to one. Nevertheless, self-influences are implicitly present to ensure conformity with Definition 1. Readers can infer an agent’s self-influence by subtracting the sum of its visible incoming influences from one.
(b) Each plot shows agents’ state evolution over time.
Figure 2: Examples of Dissensus in $\mbox{SOM}^{-}$ Models. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Initial state vector: $\mathbf{B}^{0}=(1.0,0.8,0.5,0.2,0.0)$; tolerance radii $\tau=(0.2,0.15,0.1,0.15,0.2)$; majority thresholds equal to $0.5$ for each agent.
In the example shown in Fig. 2, Agent 3 serves as the sole bridge connecting two distinct graph components: {Agents 1, 2} and {Agents 4, 5}. Due to the tolerance radius and majority threshold parameters, Agent 3 becomes effectively isolated due to its central position. Specifically, Agent 3’s belief ($0.5$) lies outside the tolerance radius of agents in both components, causing it to have no neighbors with sufficiently similar opinions. Consequently, Agent 3 becomes perpetually silent after the first round.
Meanwhile, the opposing influences from both components on Agent 3’s belief update equation cancel each other out the “pull” from the left component (Agents 1, 2) is balanced by an equal and opposite “pull” from the right component (Agents 4, 5). This results in Agent 3’s belief remaining fixed while it stays silent. In contrast, agents within each component maintain beliefs within each other’s tolerance radii (after the first round), ensuring they continue to speak and influence one another. As a result, each component converges to its own local consensus, while the perpetually silent Agent 3 can no longer transmit influence between them. The graph thus becomes effectively disconnected, preventing global consensus despite the original strong connectivity.
3.4 Dissensus in $\mbox{SOM}^{+}$ Models
We now prove through an insightful counterexample that, unlike in the memoryless $\mbox{SOM}^{-}$ models, consensus is not guaranteed in $\mbox{SOM}^{+}$ models for clique graphs. This negative result shows that incorporating a minimal yet natural memory mechanism (where only the most recent public opinions are considered) can actually hinder consensus.
While $\mbox{SOM}^{+}$ models share similarities with $\mbox{SOM}^{-}$ ones regarding perpetual silence, it introduces a unique phenomenon where the entire graph can enter and indefinitely remain in a silent state. This distinction stems from how silent agents influence the opinion update in $\mbox{SOM}^{+}$.
Recall from Section 2.3.2 that $\mathtt{pub}\mathbf{B}^{t}$ is the public state of opinion at time $t$ and represents the most recent public opinion of each agent. Henceforth, we will refer to $\mathtt{pub}\mathbf{B}^{t}_{i}$ as the public opinion of agent $i$. In $\mbox{SOM}^{+}$, speaking agents influence the opinions of their neighbors as in $\mbox{SOM}^{-}$. Nevertheless, unlike in $\mbox{SOM}^{-}$, silent agents influence their neighbors with their public opinion.
Hence, if agent $i$’s opinion converges to a value where most recent public opinions of more than a fraction $\mathcal{M}_{i}$ of its neighbors fall outside its tolerance range, the agent will remain perpetually silent. This leads to scenarios where agents withdraw from discourse, while other agents continue to be influenced by these outdated public opinions.
Unlike $\mbox{SOM}^{-}$, where opinions of silent agents are disregarded, $\mbox{SOM}^{+}$ allows for the persistence of unchanged public opinions indefinitely. This can result in dissensus (i.e., lack of convergence to a consensus) due to the formation of public opinions that no longer reflect the current opinions of silent agents.
3.4.1 Counter-example to Consensus.
Consider the $\mbox{SOM}^{+}$ model $M=(G,\mathbf{B}^{0},\mathbf{S}^{0},\mu_{G})$ with $G$ as the Clique in Fig.3(a) and $\mathbf{B}^{0}=(1.0,0.9,0.1,0.0)$, tolerance radius $\tau_{i}=0.1$, majority threshold $\mathcal{M}_{i}=0.5$ for each agent $i$ in $G$. All agents initially have more than half of their neighbors’ opinions outside their tolerance radius. Consequently, all agents become silent at $t=1$. At this point, each agent’s updated opinion places all other agents’ opinions outside its tolerance radius. The agents’ opinions then converge to distinct values, determined by the initial opinions $\mathbf{B}^{0}$, which perpetuates the condition for silence. This state persists indefinitely, as the convergence values maintain the silence condition for all agents. This results in dissensus as illustrated in Fig.3(b) .
The above counter-example demonstrates how the $\mbox{SOM}^{+}$ models can lead to complete silence and opinion divergence in a clique, preventing the possibility of consensus.
(a) Clique influence graph (self influence excluded)
(b) Each plot shows agents’ state evolution over time.
Figure 3: Counter-Example to Consensus in an $\mbox{SOM}^{+}$ Model. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Initial state vector: $\mathbf{B}^{0}=(1.0,0.9,0.1,0.0)$, tolerance radius and majority threshold equal to $0.1$ and $0.5$ for each agent.
4 Experimenting with the Spiral of Silence
This section demonstrates how $\mbox{SOM}^{-}$ and $\mbox{SOM}^{+}$ models can capture key dynamics of the spiral of silence through agent-based simulations. We first showcase an $\mbox{SOM}^{-}$ model that reinforces dominant opinions, enabling vocal minorities to disproportionately influence public discourse despite numerical inferiority. Next, we illustrate an $\mbox{SOM}^{+}$ model revealing an apparent dichotomy: agents privately converge to consensus while maintaining divergent public stances, mirroring social environments where outdated expressions mask underlying agreement [18].
To validate scalability, we simulate networks ranging from 4 to over 2.1 million agents ($2^{2}$ to $2^{21}$) with varying connectivity, leveraging a modified version of Ross et al.’s algorithm [24] to generate networks with power-law ("rich-get-richer") degree distributions and small-world properties. Small-world properties are characterized by short average path lengths between nodes while maintaining high local clustering. This reflects real social networks’ structure where information spreads efficiently despite limited local connections, making our simulations more realistic for studying opinion dynamics. Indeed, studies have shown social networks exhibit 6 degrees of separation [28], that is, any one person is on average 6 connections away from any other.
4.1 Reinforcement of Dominant Views
Figure 4 illustrates how $\mbox{SOM}^{-}$ models capture the essence of the Spiral of Silence theory. In this strongly connected graph, a vocal minority effectively dominates the discourse, causing the silent majority to converge toward the perceived majority opinion.
The network has two groups: Group 1 (Agents 1, 2, 3, 6, 7, and 8) with opinions in the lower half of the spectrum, and Group 2 (Agents 4 and 5) with opposing views near the higher end. Despite Group 1 being the actual majority, the network topology and differing tolerance radii $\tau$ lead to Group 2 dominating the opinion dynamics. Group 1 agents lack intra-group connections but are all linked to Group 2 agents. Also, Group 2’s significantly larger tolerance radii allow them to remain non-silent more frequently.
(b) Each plot shows agents’ state evolution over time.
Figure 4: Silent Majority vs. Vocal Minority: Opinion Dynamics in $\mbox{SOM}^{-}$. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Here $\mathbf{B}^{0}=(0.1,0.2,0.15,1.0,0.85,0.25,0.3,0.05)$ and $\tau=(0.1,0.05,0.1,0.85,0.6,0.05,0.1,0.05)$ and majority thresholds are set to $0.5$.
This configuration results in the vocal minority (Group 2) disproportionately influencing the network. The silent majority (Group 1) remains quiet for most of the update process, only becoming active when opinions have already shifted closer to the perceived majority view. This example demonstrates how the $\mbox{SOM}^{-}$ models can simulate scenarios where a minority opinion, through strategic positioning and persistent vocalization, can shape the overall opinion landscape, even when numerically outnumbered.
4.2 Hidden Consensus in $\mbox{SOM}^{+}$ Models
$\mbox{SOM}^{+}$ models reveal a noteworthy phenomenon: the possibility of reaching a consensus that remains undetected by the agents themselves. Figure 5 illustrates this scenario using a clique graph with four agents.
Initially, the agents hold diverse opinions ($\mathbf{B}^{0}=(1.0,0.9,0.1,0.0)$). However, due to the graph’s influence structure, all agents become silent after $t=0$. Despite this silence, their private opinions converge to a common value over time. Crucially, this convergence occurs without any further public expression of opinions, leaving each agent unaware of the emerging consensus. This hidden consensus phenomenon mirrors real-world scenarios in social media where individuals may unknowingly share common views while perceiving disagreement due to outdated public expressions [18].
(a) Influence graph self influence excluded
(b) Each plot shows agents’ state evolution over time.
Figure 5: Silent Consensus Scenario. Triangles represent silent agents, circles non-silent ones. Colored areas indicate opinion values within each agent’s tolerance radius. Initial state vector: $\mathbf{B}^{0}=(1.0,0.9,0.1,0.0)$; tolerance radii $\tau=(0.2,0.1,0.05,0.15)$; majority thresholds equal to $0.5$ for each agent.
4.3 Scaling to Large-Scale Network Simulations
While small-scale simulations illustrate core mathematical behaviors, they do not capture the complexity of real-world social networks. To bridge this gap, we developed a scalable simulation platform supporting networks of over 2.1 million agents ($2^{21}$), enabling analysis of emergent phenomena in realistic settings.
Our implementation combines a modified preferential attachment algorithm [1] to generate networks with small-world properties [29] and power-law (rich-get-richer) degree distributions. To address computational intensity, we employ parallel processing via Scala and Akka Actors [20, 19], with results persisted in PostgreSQL for efficient querying. For each experiment, we generated networks with density (minimum neighbor count) ranging from 1 to 15, executing 1,024 simulations per parameter combination. Consensus rates were then calculated across all runs for statistical significance. All agents have their tolerance radius and majority threshold set to $0.1$ and $0.5$, respectively, and their initial beliefs as random uniformly distributed values from 0 to 1. The code is available here.
4.3.1 Contrasting Silence Opinion Models.
The $\mbox{SOM}^{+}$ and $\mbox{SOM}^{-}$ models exhibit dramatically opposing behaviors in consensus formation as seen in Fig.6. For $\mbox{SOM}^{+}$ models, consensus becomes increasingly rare beyond 512 agents ($2^{9}$), occurring in only 0.097% of simulations (1/1024 runs) at this threshold, exclusively for density 4. Beyond this scale, no simulations reached consensus. This aligns with social media dynamics where persistent disagreement emerges, as agents’ historical opinions create perception gaps between private consensus and public expression [18]. Even a minimal notion of memory (retaining only most recent public opinions) proves sufficient to sustain dissensus.
Figure 6: Heatmap indicating consensus % across different network sizes (y-axis) and densities (x-axis) for both $\mbox{SOM}^{+}$ and $\mbox{SOM}^{-}$ models. Lighter hues indicate a higher rate of consensus, while darker hues indicate a lower rate. All tolerance radii and majority thresholds are equal to $0.1$ and $0.5$ resp.
Conversely, $\mbox{SOM}^{-}$ models exhibit the opposite scalability behavior. Once a certain density threshold is reached, consensus consistently emerges in all simulations. Larger networks require slightly higher connectivity: for 2,048 agents ($2^{11}$), a density of 3 suffices, while for 2.1 million agents ($2^{21}$), a density of 4 is needed. This highlights how the spiral of silence depends on network structure. In sparse networks, some agents may remain silent indefinitely due to critical bridge nodes (see Section 3.3). In contrast, denser networks make it increasingly unlikely for bridges to be disconnected, since disrupting consensus would require silencing more than $d$ agents (for density dd)—a rare occurrence in networks of this type.
5 Conclusions and Related Work
We extended the classical DeGroot framework to incorporate key social dynamics described by the Spiral of Silence. Our contributions highlight how the addition of memory, even in a limited form, as in $\mbox{SOM}^{+}$ models, introduces significant complexity to consensus-building processes. We demonstrated that consensus, while achievable for cliques in $\mbox{SOM}^{-}$ models, is no longer guaranteed, in contrast with DeGroot models, in arbitrary strongly-connected aperiodic graphs. This points to the impact that silence and memory can have on opinion formation in social networks. It also offers insights into the challenges of converging to consensus in real-world scenarios. We also discuss simulations reflecting predictions of the Spiral of Silence, such as the reinforcement of dominant views and hidden consensus. Finally, we discussed the simulation in our models of large-scale graphs reflecting the small-world topology of real social networks.
Future Work. Experimentation and validation on real social networks typically involve continuously collecting user opinions and inferring influence via methods such as stance detection and influencer identification from natural language processing, machine learning and graph mining. We defer this challenging task to future work.
Related Work. To our knowledge, no prior work has extended DeGroot-based models to incorporate the Spiral of Silence. Some recent studies have examined the Spiral of Silence in agent-based networks. E.g., [24] models manipulative actors in social networks using agents with fixed binary opinions (agree/disagree) who choose whether to express them based on the prevailing opinion climate. In [8], the authors explore how the number of communities and their connectivity influence perceived opinion climates. However, these studies do not address opinion updates, convergence to consensus, or memory, which are central aspects of this paper.
In [12, 33] the stabilization property was studied in majority-based opinion models, where individuals adopt the most prevalent opinion in their social circles (their neighbors). Although these works focus on majority-based opinion dynamics, similar to models inspired by Spiral of Silence, their models are exclusively memoryless with only two-value discrete opinions.
Recent studies [7, 25, 32] explored conflict, disagreement, and polarization using the Friedkin and Johnsen (FJ) model. This model extended DeGroot’s by adding a fixed internal opinion and modeling stubbornness, where agents’ expressed opinions reflected their internal beliefs. While silent agents in $\mbox{SOM}^{+}$ and stubborn agents in the FJ model shared some similarities, their behavior differed. In $\mbox{SOM}^{+}$, silent agents could later become active, allowing their public opinions to change. In contrast, stubborn agents in the FJ model did not shift in this way. Thus, although neither model guaranteed consensus, their opinion dynamics diverged.
The exclusion of silent agents in $\mbox{SOM}^{-}$ models amounts to having edges (influences) disappearing and reappearing during opinion evolution which reflects the dynamic influence nature of this model. There are several works studying dynamic influence in opinion formation. The work [15] introduces a version of the DeGroot model in which self-influence changes over time while the influence on others remains the same. The works [10, 11] explore convergence and stability, respectively, in models where influences change over time. The work [5] demonstrates how asynchronous communication, when combined with dynamic influence, can prevent consensus. None of this work deals with the dynamics derived from the Spiral of Silence.
Appendix 0.A Appendix
See 1
@@ Appendix 0.A Appendix
See 2
@@
$\mathbf{B}^{t+1}_{i}=(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot\mathbf{B}^{t}_{i}+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot\mathbf{B}^{t}_{j}$
Now, as $\mathbf{B}^{t}_{k}\leq max(\mathbf{B}^{t})$ for all $k$ $\in$ AA, we can conclude that:
$\mathbf{B}^{t+1}_{i}\leq(1-\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j})\cdot max(\mathbf{B}^{t})+\sum_{j\in N_{i}}I_{ji}\cdot\mathbf{S}^{t}_{j}\cdot max(\mathbf{B}^{t})=max(\mathbf{B}^{t})$
Thus, $\mathbf{B}^{t+1}_{i}\leq max(\mathbf{B}^{t})$ as wanted. The proof that $min(\mathbf{B}^{t})\leq\mathbf{B}^{t+1}_{i}$ is similar.
See 3